# Hierarchical Selective Classification

### v1

@@ 5 Experiments
Table 2: Results (mean scores) comparing the hierarchical selective threshold algorithm (Algorithm 2) with DARTS, repeated 1000 times for each model and target accuracy with a randomly drawn calibration set of 5,000 samples, applied to 1115 models trained on ImageNet1k (meaning we evaluated each algorithm 1,115,000 times). $1-\delta$ is set at 0.9.
@@
(2) Pretraining on larger ImageNet datasets benefits hierarchical selective performance, with certain models achieving up to a 40% improvement. However, the improvement rates vary significantly: not all models experience the same benefit from pretraining, and some may experience minimal or no improvement. While semi-supervised learning also shows improvement in hierarchical selective performance, the relatively low number of models tested (only 11) makes it challenging to draw firm conclusions.
@@
(4) Linear-probe CLIP significantly outperforms zero-shot CLIP in HSC: We compared pairs of models with identical backbones, where one is the original zero-shot model, and the other was equipped with a linear-probe, that is, it uses the same frozen feature extractor but has an added head trained to classify ImageNet1k. The zero-shot models evaluated are the publicly available models released by OpenAI. The mean relative improvement from zero-shot CLIP to linear-probe CLIP is 45%, with improvement rates ranging from 32% to 53%. We hypothesize that the hierarchical selective performance boost may be related to better calibration or ranking abilities. Specifically, we observe that all CLIP models equipped with linear-probes had significantly higher AUROC than their zero-shot counterparts, indicating superior ranking. Figure 7 in Appendix I shows these results in more detail.
@@ 6 Related Work
In selective classification, several alternative methods for confidence thresholding have been developed [5, 27]. However, these methods generally necessitate some form of training. This work is focused on post-hoc methods that are compatible with any pretrained classifier, which is particularly advantageous for practitioners. Furthermore, despite the availability of other options, threshold-based rejection even with the popular Softmax confidence score continues to be widely used [16, 19], and can be readily enhanced with temperature scaling or other techniques [6, 14, 16].
@@
Various works leverage a hierarchy of classes to improve leaf predictions of flat classifiers [45, 4, 24] but comparatively fewer works explore hierarchical classifiers. [45] optimized a “win” metric comprising a weighted combination of likelihoods on the path from the root to the leaf. [4] focused on reducing leaf mistake severity, measured by graph distance. They introduced a hierarchical cross-entropy loss, as well as a soft label loss that generalizes label smoothing. [24] claimed that both methods in [4] result in poorly calibrated models, and proposed an alternative inference method. [11] proposed the Dual Accuracy Reward Trade-off Search (DARTS) algorithm, which attempted to obtain the most specific classifier for a user-specified accuracy constraint. They used information gain and hierarchical accuracy as two competing factors, which are integrated into a generalized Lagrange function to effectively obtain multi-granularity decisions. Additional approaches include the Level Selector network, trained by self-supervision to predict the appropriate level in the hierarchy [23]. [40] proposed a loss based on [46] and performed inference using a threshold-based inference rule. Other works allow non-mandatory leaf node prediction, although not directly addressing the accuracy-specificity trade-off [32, 46]. The evaluation of hierarchical classifiers has received relatively little attention previous to our research. The performance profile of a classifier could be inferred from either the average value of a metric across the operating range or by observing operating curves that compare correctness and exactness or information precision and recall [40], or measuring information gain [11, 46]. Set-valued prediction [20] tackles a similar problem to hierarchical classification by allowing a classifier to predict a set of classes. Although both approaches handle uncertainty by predicting a set of classes, the HSC framework hard-codes the sets as nodes in the hierarchy. This way, HSC injects additional world knowledge contained in the hierarchical relations between the classes, yielding a more interpretable prediction. Conformal risk control [3] is of particular interest because it constrains the prediction sets to be hierarchical. Appendix H shows a comparison of this baseline to our method.
@@ 7 Concluding Remarks
This paper presents HSC, an extension of selective classification to a hierarchical setting, allowing models to reduce the information in predictions by retreating to internal nodes in the hierarchy when faced with uncertainty. The key contributions of this work include the formalization of hierarchical risk and coverage, the introduction of hierarchical risk-coverage curves, the development of hierarchical selective inference rules, and an efficient algorithm to find the optimal confidence threshold for a given target accuracy. Extensive empirical studies on over a thousand ImageNet classifiers demonstrate the advantages of the proposed algorithms over existing baselines and reveal new findings on training methods that boost hierarchical selective performance.
However, there are a few aspects of this work that present opportunities for further investigation and improvement: (1) Our approach utilizes softmax-based confidence scores. exploring alternative confidence functions and their impact on hierarchical selective classification could provide further insights; (2) While we have identified certain training methods that boost hierarchical selective performance, the training regimes were not optimized for hierarchical selective classification. Future research could focus on optimizing selective hierarchical performance. (3) Although our threshold algorithm is effective, it could be beneficial to train models to guarantee specific risk or coverage constraints supplied by users, in essence constructing a hierarchical “SelectiveNet” [18];
@@
Appendix A Max-Coverage Inference Rule Algorithm
@@
Appendix B Jumping Inference Rule
@@
Appendix C Hierarchy Traversal for Threshold Finding
@@
Appendix D Proof of Theorem 1
@@
In conformal prediction, the probability $P(\theta_{n+1}\leq\hat{\theta})$ is called “marginal coverage” (not to be confused with the hierarchical selective definition of coverage in section LABEL:sec:hsc). The analytic form of marginal coverage given a fixed calibration set, for a sufficiently large test set, is shown by [41] to be
@@
Appendix E Comparison of Inference Rules Without Temperature Scaling
Table 3: hAURC and gain results for each inference rule. Mean results of 1115 pretrained ImageNet1k models.
@@ Appendix E Comparison of Inference Rules Without Temperature Scaling
Appendix F Hierarchical Selective Classification improves Confidence Calibration
Confidence calibration refers to the problem of predicting probability estimates representative of the true correctness likelihood. The calibration of both hierarchical and non-hierarchical versions of selective classification is a highly intriguing topic that has not been thoroughly explored in previous research. Fisch et al. [15] suggested that often, but not always, more confident predictions can also be more calibrated.
We measure calibration using the Expected Calibration Error (ECE) [30], a commonly used metric for calibration. For each model, we computed the ECE for each hierarchical coverage. We then computed the area under the ECE-Coverage curve. Subsequently, we calculated the area under the ECE-Coverage curve, and compared these areas similarly to the hAURC metric described in section 5.
Our analysis, which includes results from 1,115 models, reveals that selective classification not only improves risk, but also improves calibration. Additionally, the choice of hierarchical inference rule significantly impacts whether calibration improves or degrades. On average, across the 1,115 models, the mean area under the ECE-Coverage curve when applying the selective inference rule was 36.58. For Max-Coverage [40], the mean area was 55.89, while for the climbing inference rule, the mean area was the lowest at 12.44 (a smaller area indicates better calibration, areas were multiplied by 1000). Interestingly, unlike hAURC, hierarchical inference rules do not necessarily show improvement over the selective baseline: the area for Max-Coverage is significantly higher than that for the selective rule. Figure 5 compares two ECE-Coverage curves for the same model, demonstrating that the climbing inference rule achieves a better curve compared to the selective baseline.
Figure 5: Comparison of selective and climbing hierarchical ECE-Coverage curves rules with ViT-L/16-384 as the base classifier. This example represents the prevalent case, in which climbing outperforms the “hierarchically-ignorant” selective inference rule.
@@ Appendix H Conformal Risk Control Results
Figure 6: Mean accuracy error comparison between the hierarchical selective threshold algorithm (Algorithm 2), conformal risk control (CRC) and DARTS, repeated 1000 times for each model and target accuracy with a randomly drawn calibration set of 5,000 samples, applied to 1115 models trained on ImageNet1k (meaning we evaluated each algorithm 1,115,000 times). $1-\delta$ is set at 0.9.
@@ Appendix I CLIP Zero Shot VS Linear Probe
Appendix J Technical Details
@@
The evaluation was performed on 1115 vision models pretrained on ImageNet1k [10], and 6 models pretrained on iNat-21 [22] (available in timm 0.9.16 [43] and torchvision 0.15.1 [28]).
All experiments were conducted on a single machine with one Nvidia A4000 GPU. The evaluation of all experiments on a single GPU took approximately one week.

### v2

@@ 5 Experiments
Table 2: Results (mean scores) comparing the hierarchical selective threshold algorithm (Algorithm 2) with DARTS, repeated 1000 times for each model and target accuracy with a randomly drawn calibration set of 5,000 samples, applied to 1,115 models trained on ImageNet1k (meaning we evaluated each algorithm 1,115,000 times). $1-\delta$ is set at 0.9.
@@
(2) Pretraining on larger ImageNet datasets benefits hierarchical selective performance, with certain models achieving up to a 40% improvement. However, the improvement rates vary significantly: not all models experience the same benefit from pretraining. Semi-supervised learning also shows a noticeable improvement in hierarchical selective performance.
@@
(4) Linear-probe CLIP significantly outperforms zero-shot CLIP in HSC: We compared pairs of models with identical backbones, where one is the original zero-shot model, and the other was equipped with a linear-probe, that is, it uses the same frozen feature extractor but has an added head trained to classify ImageNet1k. The zero-shot models evaluated are the publicly available models released by OpenAI. The mean relative improvement from zero-shot CLIP to linear-probe CLIP is 45%, with improvement rates ranging from 32% to 53%. We hypothesize that the hierarchical selective performance boost may be related to better calibration or ranking abilities. Specifically, all CLIP models with linear-probes showed significantly higher AUROC than their zero-shot counterparts, indicating superior ranking. Figure 7 in Appendix I shows these results in more detail.
Figure 5: Aggregated (mean and SEM) CC curves of 1,115 ImageNet models.
Since hAURC is closely related to accuracy, we sought to determine whether its improvement could be attributed solely to gains in predictive performance from the training regimes. Our analysis, detailed in Appendix J, suggests this is not the case. Additionally, we explored a training regime specifically aimed at enhancing HSC performance. Our approach is described in Appendix K.
HSC Improves Confidence Calibration: We plotted the Calibration-Coverage (CC) Curves (defined in Section 2 for 1,115 ImageNet models and compared the curves produced by the Selective inference rule to those from Climbing. Figure 5 presents the aggregated CC curve for all 1,115 models across the two inference rules. The results indicate that HSC not only improves risk, but also improves calibration, with the Climbing inference rule nearly always outperforming Selective.
@@ 6 Related Work
In selective classification, several alternative methods for confidence thresholding have been developed [5, 31, 38]. However, these methods generally necessitate some form of training. This work is focused on post-hoc methods that are compatible with any pretrained classifier, which is particularly advantageous for practitioners. Furthermore, despite the availability of other options, threshold-based rejection even with the popular Softmax confidence score continues to be widely used [18, 21], and can be readily enhanced with temperature scaling or other techniques [6, 15, 18].
@@
Various works leverage a hierarchy of classes to improve leaf predictions of flat classifiers [51, 4, 26] but comparatively fewer works explore hierarchical classifiers. [51] optimized a “win” metric comprising a weighted combination of likelihoods on the path from the root to the leaf. [4] focused on reducing leaf mistake severity, measured by graph distance. They introduced a hierarchical cross-entropy loss, as well as a soft label loss that generalizes label smoothing. [26] claimed that both methods in [4] result in poorly calibrated models, and proposed an alternative inference method. [12] proposed the Dual Accuracy Reward Trade-off Search (DARTS) algorithm, which attempted to obtain the most specific classifier for a user-specified accuracy constraint. They used information gain and hierarchical accuracy as two competing factors, which are integrated into a generalized Lagrange function to effectively obtain multi-granularity decisions. Additional approaches include the Level Selector network, trained by self-supervision to predict the appropriate level in the hierarchy [25]. [46] proposed a loss based on [52] and performed inference using a threshold-based inference rule. Other works allow non-mandatory leaf node prediction, although not directly addressing the accuracy-specificity trade-off [36, 52]. The evaluation of hierarchical classifiers has received relatively little attention previous to our research. The performance profile of a classifier could be inferred from either the average value of a metric across the operating range or by observing operating curves that compare correctness and exactness or information precision and recall [46], or measuring information gain [12, 52]. Set-valued prediction [22] and hierarchical multi-label classification [45] tackle a similar problem to hierarchical classification by allowing a classifier to predict a set of classes. Although both approaches handle uncertainty by predicting a set of classes, the HSC framework hard-codes the sets as nodes in the hierarchy. This way, HSC injects additional world knowledge contained in the hierarchical relations between the classes, yielding a more interpretable prediction. Conformal risk control [3] is of particular interest because it constrains the prediction sets to be hierarchical. Appendix H shows a comparison of this baseline to our method.
Calibration is another important aspect of uncertainty estimation [27, 58, 18, 16]. Fisch et al. [16] demonstrate that selectively abstaining from uncertain predictions may improve calibration.
@@ 7 Concluding Remarks
This paper presents HSC, an extension of selective classification to a hierarchical setting, allowing models to reduce the information in predictions by retreating to internal nodes in the hierarchy when faced with uncertainty. The key contributions of this work include the formalization of hierarchical risk and coverage, hierarchical calibration, the introduction of hierarchical risk-coverage curves and hierarchical calibration-coverage curves, the development of hierarchical selective inference rules, and an efficient algorithm to find the optimal confidence threshold for a given target accuracy. Extensive empirical studies on over a thousand ImageNet classifiers demonstrate the advantages of the proposed algorithms over existing baselines and reveal new findings on training methods that boost hierarchical selective performance. However, there are a few aspects of this work that present opportunities for further investigation and improvement: (1) Our approach utilizes softmax-based confidence scores. exploring alternative confidence functions and their impact on hierarchical selective classification could provide further insights; (2) While we have identified certain training methods that boost hierarchical selective performance, the training regimes were not optimized for hierarchical selective classification. Future research could focus on optimizing selective hierarchical performance. (3) Although our threshold algorithm is effective, it could be beneficial to train models to guarantee specific risk or coverage constraints supplied by users, in essence constructing a hierarchical “SelectiveNet” [20].
8 Acknowledgments
The research was partially supported by Israel Science Foundation, grant No 765/23
@@ 8 Acknowledgments
Appendix A Hierarchical Severity Risk
The presence of a hierarchy allows us to measure the severity of incorrect predictions, instead of treating all mistakes as equal. For instance, classifying a Labrador as a Golden Retriever is a much milder mistake compared to classifying it as a snake. We use the coverage of the Lowest Common Ancestor (LCA) of the predicted node $\hat{v}$ and the label v to define a novel loss function that accounts for mistake severity: $l^{H}(\hat{v},y)=1-\frac{\phi(LCA(\hat{v},v))}{\phi(\hat{v})}.$
Building upon the above example, the LCA of Labrador Retriever and Golden Retriever is the node ’Retriever’, which has significantly higher coverage relative to the prediction ’Golden Retriever’, in contrast to ’Vertebrate’, the LCA of Labrador Retriever and Snake, resulting in the former misclassification having lower risk compared to the latter.
We provide below the results of our experiments on ImageNet, complementing Table 1:
Table 3: Comparison of mean hAURC and hierarchical gain results for the selective, Max-Coverage (MC) and Climbing inference rules applied to 1,115 models trained on ImageNet1k, using the hierarchical severity risk.
ImageNet-1k (1,115 models)
Inference Rule hAURC Hier. Gain hAURC Hier. Gain
($\times$ 1000) (%) ($\times$ 1000) (%)
Selective 24.98$\pm$0.29
MC 24.50$\pm$0.34 3.31$\pm$0.30
Climbing 24.04$\pm$0.31 4.71$\pm$0.22
The results align with those presented in the main body of the paper, where the risk is the the 0/1 loss. Climbing remains the inference rule with the best (lowest) hAURC and the highest hierarchical gain. In practice, the results of both losses are highly correlated.
Appendix B Max-Coverage Inference Rule Algorithm
@@
Appendix C Jumping Inference Rule
@@
Appendix D Hierarchy Traversal for Threshold Finding
@@
Appendix E Proof of Theorem 1
@@
In conformal prediction, the probability $P(\theta_{n+1}\leq\hat{\theta})$ is called “marginal coverage” (not to be confused with the hierarchical selective definition of coverage in section 2). The analytic form of marginal coverage given a fixed calibration set, for a sufficiently large test set, is shown by [47] to be
@@
Appendix F Comparison of Inference Rules Without Temperature Scaling
Table 4: hAURC and gain results for each inference rule. Mean results of 1,115 pretrained ImageNet1k models.
@@ Appendix H Conformal Risk Control Results
Figure 6: Mean accuracy error comparison between the hierarchical selective threshold algorithm (Algorithm 2), conformal risk control (CRC) and DARTS, repeated 1000 times for each model and target accuracy with a randomly drawn calibration set of 5,000 samples, applied to 1,115 models trained on ImageNet1k (meaning we evaluated each algorithm 1,115,000 times). $1-\delta$ is set at 0.9.
@@ Appendix I CLIP Zero Shot VS Linear Probe
Appendix J Training Regimes Improvement in hAURC VS Accuracy
In Table 6 we compare the relative improvement in hAURC to the relative improvement in accuracy. It can be seen that although the training regimes with the highest and lowest improvement in hAURC have the highest and lowest improvement in accuracy, the improvement in hAURC does not determine the rate of improvement in accuracy.
Table 6: Relative improvement in Accuracy compared to relative improvement in hAURC.
Training Regime Median Accuracy Improvement (%) Median hAURC Improvement (%)
CLIP 7.90 38.14
Pretraining 1.96 19.47
SSL 4.12 18.97
Distillation 1.48 11.73
Adversarial 0.85 5.92
Appendix K Training Regime to Optimize HSC Performance
Since hAURC cannot be optimized directly, we developed an alternative that could be optimized. Our best-performing method entailed training models to predict the lowest common ancestor (LCA) of pairs of samples, including identical pairs (intended for testing). To achieve this, we fine-tuned models using the hierarchical loss we proposed in Appendix A.
All models were fine-tuned on the ImageNet validation set for 20 epochs using the SGD optimizer with a warmup scheduler and a batch size of 2048 on a single NVIDIA A40 GPU. After fine-tuning models of various architectures (including ResNet50, ViTs, EfficientNet, and others) with this loss function and utilizing various other configurations, such as leveraging in-batch negatives for a richer training signal and multiple classification heads to prevent deterioration in the model’s accuracy, the improvement in hAURC we observed ranged from 3% to 5% compared to the pretrained baseline. We encourage future work to explore this direction further.
Appendix L Technical Details
@@
The evaluation was performed on 1,115 vision models pretrained on ImageNet1k [11], and 6 models pretrained on iNat-21 [24] (available in timm 0.9.16 [49] and torchvision 0.15.1 [32]), both available under the MIT license.
All experiments were conducted on a single machine with one Nvidia A4000 GPU. The evaluation of all experiments on a single GPU took approximately two weeks.